OpenStudy (anonymous):

If the sides of a square are lengthed by 6cm the area becomes 225cm^2. Find the length of the side of the original square?

OpenStudy (anonymous):

(x+6)(x+6)=225 x^2+12x+36=225 x^2+12x-189=0 x=9, -21 -21 is a negative length so the answer is 9